OpenStudy (anonymous):

Consider the function f(x)=x^2+1. List the first five approximations (i.e. through x5) Newton’s Method gives for this function using the initial guess x1=2. Does Newton’s Method appear to be converging? If not, what appears to be happening?

OpenStudy (amistre64):

newtons method is simply to use equation of the tangent line: newX = f'(x) -f'(x)(x) + f(x)

OpenStudy (amistre64):

f'(x) = 2x in this case; so... newX = 2(2) -2(2) + 5 ;

OpenStudy (amistre64):

that seems a bit off...

OpenStudy (amistre64):

slope = f'(2) = 2(2) = 4 , at point(2,5) newX = 4x -4(2) + 5 = 0 = 4x -3 = 0 = 3/4

OpenStudy (amistre64):

t{n} = 2(t{n-1})x -2(t{n-1}) + (t{n-1})^2 + 1 ; looks a bit convoluted, but I think its the recurrsion for this :)